# Inside you are many wolves: Using cognitive models to interpret value trade-offs in LLMs

### v1

@@ 5.1 Experimental vignettes
We provide models with the same set of vignettes given to human participants in Yoon et al., [92], which describe socially sensitive situations in which a speaker must convey their judgement of a listener’s creation (e.g. a poem, presentation, cake, etc.). The speaker’s actual opinion, or true state $s$, is expressed on a scale from 1 to 5 stars, where 1 is the lowest or most negative opinion, and 5 is the highest. [^3] We deviate from the original paper’s 0-3 heart scale to provide LLMs with a scale that is most natural to their training data, particularly online reviews. We find that this 1-5 star scale captures the semantic range of the available utterance options better than the original 0-3 scale. We present models with the set of eight utterance options $u$ (four descriptor words and their negations) in a multiple choice format:
@@
Manipulations of vignette framing
Since LLMs are increasingly being used to take on diverse roles, such assistants to users and agents acting in their own capacity, we consider how these points of view might affect the values an LLM prioritizes. To assess this, we extend the original third-person framing of the above scenario (simulating an LLM-as-judge) to also evaluate LLMs on the first- and second-person framings of these vignettes (see Section A.3 for model prompts).
LLM response evaluation
To control for ordering effects, utterance options were presented to the models in a random order. The majority of models’ generations adhered to the specified multiple-choice format, but in cases where they did not, we used GPT-4o as a judge to determine the LLM’s answer choice (see Section A.2). Then, among the valid responses, LLMs’ choice of utterance for a given scenario and true state (e.g. a poem that was worthy of 4 stars) was measured as the normalized probabilities assigned to each possible utterance option (see Section C.1 for response distributions).
@@ 5.2 Inferring cognitive model parameters
We implemented the inference model in Stan [9], a probabilistic programming language, and used the default Hamiltonian Monte Carlo implemented in Stan (No-U-Turn sampler, Hoffman et al., [30])) to perform approximate inference of model parameters. We ran 4 chains, with 2000 warm-ups and 2000 samples for each chain. For the results, we report the posterior mean as well as the 95% high density interval of the inferred parameters $\Theta$ fitted on the transformed LLM utterance preference data $\mathcal{M}$ (see Section B.2 for further implementation details).
@@
To infer our desired cognitive model parameters $\boldsymbol{\omega}$ and $\phi$, we require an estimate of the parameter $\theta$, the probability that the utterance $u$ is true of state $s$. To obtain this, we query LLMs with a modified version of the main task where the following question is appended to the above Scenario, in its original third-person framing:
@@
For both open- and closed- source LLMs, we measure the model’s “endorsement” of a particular utterance $u$ for state $s$ as the posterior mean of the probability of success (i.e. a “yes” response for $u$ describing $s$) under a Beta-Binomial model with a uniform prior following [92]. We obtain a total of 52 samples (4 random combinations of speaker and listener names for each creation $c$) per state-utterance pair, replicating the human study sample size $(n=51)$ (see Section C.2 for an example of LLMs’ responses on this sub-task).
Figure 2: Closed-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$ (purple), and projected mixture of informational and social utilities $\phi$ (magenta), according to the cognitive model for LLMs with varying degrees of reasoning ’effort’. Error bars indicate 95% high density region averaged across results from three framing manipulations. Model parameters estimated from human behavioral data for different goal conditions are taken from Yoon et al., [92].
@@ 6.1 Human baseline
In the original study [92] , human participants were asked to assume the role of the speaker, and to choose an utterance according to one of three goal conditions: trying to be informative, trying to be social (i.e. kind), or both. The work finds that speakers who have the conflicting goals of being both informative and kind will use more indirect speech when describing a bad state (e.g. they describe a cake that deserves only 1 star as ‘not amazing’). This behavior serves to “save face” (i.e. optimize presentational and social utilities), while still conveying useful information about the true state. It suggests that humans do not eschew one of their goals to increase utility along a single dimension, but rather, choose the utterances that will jointly maximize their competing utilities.
These qualitative patterns are reflected in the maximum a posteriori (MAP) estimates of the $\phi$ and $\boldsymbol{\omega}$ parameters of the $S_{2}$ model (hatched bar group in Figure 2). The relative parameter values in each goal condition provide baselines against which we can interpret a model’s default (non-goal-conditioned response). Speakers in the ‘informative’ goal condition project a balanced, but more information-leaning weighting of information and social utilities (ϕ=italic-$\phi=$0.49) than those in the social goal or combined goal conditions (0.37 and 0.36, respectively). The relative weightings of information and social utility in $S_{2}$, $\omega_{\text{inf}}$ and $\omega_{\text{soc}}$, track with these goal conditions, while humans’ $\omega_{\text{pre}}$, their value for communicating their $\phi$ to a listener, is highest for the informative goal condition (0.62), followed by the combined condition (0.54), and finally the social condition (0.44).
In the following sections, we report the results of fitting the LLMs’ responses to the second-order speaker model to obtain estimates of $\phi$ and $\boldsymbol{\omega}$, aggregated over the three manipulations of vignette framings. For disaggregated response distributions and sample results of the intermediate, first-order speaker model, see Appendix C.
Figure 3: Open-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$ (purple), and projected mixture of informational and social utilities $\phi$ (magenta), according to a cognitive model for LLMs’ training checkpoints across the RLHF process. Line variants indicate different combinations of base model and feedback dataset; rows = alignment method. Error bars indicate 95% high density region averaged across results from three framing manipulations.
@@ 6.2 Closed-source model suite
Figure 2 shows the results of fitting the reasoning and non-reasoning variants of Anthropic, Gemini, and OpenAI’s language models responses to the second-order speaker model.
@@
We begin with the inferred parameter values of $\boldsymbol{\omega}$, which measures the weightings of informational, social, and presentational utilities used by the second-order pragmatic speaker. Within the Anthropic model family, Claude-Sonnet-3.5 (no reasoning), shows a significantly lower weighting of informational utility $\omega_{\text{inf}}$ compared to its low-reasoning counterpart, Claude-Sonnet-3.7 ($t=-5.57,p=0.009$), but significantly higher social utility $\omega_{\text{soc}}$ ($t=8.70,p=0.01$). Among the OpenAI models, a similar pattern holds with significantly lower $\omega_{\text{inf}}$ for no reasoning compared to low reasoning effort ($t=-6.44,p<0.01$), but not $\omega_{\text{soc}}$ ($p=0.09$). Conversely, the Gemini-Flash models do not show a significant difference between reasoning and non-reasoning variants for any of $\boldsymbol{\omega}$ ($p=0.42$ for $\omega_{\text{inf}}$, $p=0.24$ for $\omega_{\text{soc}}$, $p=0.88$ for $\omega_{\text{pre}}$).
[…]
@@
Finally, considering the mean speaker optimality $\alpha$, averaged over reasoning variants and vignette framings, suggests that the above described weightings of utilities do factor into the models’ choice of utterances, with all model families’ $\alpha$ being higher than 1 ($\alpha_{\text{Anthropic}}$=3.55 [3.28, 3.82]; $\alpha_{\text{Gemini}}$=6.18 [5.66, 6.70]; $\alpha_{\text{OpenAI}}$=4.84 [4.25, 5.46]).
In summary, our findings from the closed-model suite suggest that 1) LLMs’ choice of utterances in this domain is sensitive to the weightings of utilities proposed by the cognitive model, 2) there are significant effects of the presence of reasoning on $\phi$, towards higher weighting on informational utility than social (but increasing the token budget devoted to reasoning effort does not have a significant effect in this regard), and 3) within model families, both the Anthropic and OpenAI models show higher $\omega_{\text{inf}}$ as a result of reasoning, while Gemini-Flash models do not show any significant changes in the relative weightings of any of the utilities in $\boldsymbol{\omega}$ as a function of reasoning.
@@ 6.3 Open-source model suite
Figure 3 shows the training dynamics of two base open-source LLMs, Qwen2.5-7B-Instruct (lighter) and Llama-3.1-8B-Instruct (darker), aligned to the UltraFeedback (dashed line) and Anthropic HH-RLHF (solid line) datasets, via DPO (top row) and PPO (bottom row). Across the different inferred parameters, we observe a number of consistent patterns within combinations of model and dataset. Across both PPO and DPO and the two feedback datasets, Qwen-instruct shows a higher $\omega_{\text{inf}}$, but lower weighting of $\omega_{\text{pre}}$ than Llama-instruct. The differences between the models’ weighting of social utility $\omega_{\text{soc}}$ are less pronounced, but still present, with Qwen-instruct generally converging to a lower weighting of social utility than Llama-instruct. The projected weighting towards informational utility in Qwen-instruct’s $\phi$, as well as its higher $\omega_{\text{inf}}$ compared to Llama-instruct aligns with prior work highlighting Qwen’s superior performance in mathematical and reasoning tasks compared to Llama [20, 94].
Turning to the effects of feedback dataset, we find that alignment to the UltraFeedback dataset most clearly results in convergence to a higher $\omega_{\text{inf}}$ for both base LLMs, than when aligned to Anthropic’s HH-RLHF dataset. In the case of $\omega_{\text{soc}}$, these differences are more pronounced as a result of PPO alignment, but still visible in the DPO case: for both base LLMs, alignment to HH-RLHF appears to result in a higher weighting of social utility than alignment to UltraFeedback. This aligns with the stated characteristics and attributes of the respective datasets, where HH-RLHF is a human feedback dataset that emphasizes more prosocial qualities like harmlessness and helpfulness, whereas UltraFeedback is a synthetic feedback dataset that contains more diverse instruction-following preferences.
@@
In general, we see that the largest shifts in utility values across all four parameters happen within the first quarter of training, consistent with earlier findings on rapid adaptation during RL post-training in mathematical domains [96].While such prior work has emphasized the significance of the base model and its pretraining data, our use of a shared supervised fine-tuning (SFT) stage on the same preference datasets across all models may attenuate these differences. Moreover, the relatively minor distinction between PPO and DPO in our results may be partly due to training both methods for only a single epoch, and the fact that the Armo-RM reward model used in PPO, was trained on subsets of the same UltraFeedback and Anthropic HH-RLHF datasets, further reducing divergence between the two approaches.
@@ 7 Discussion
Through a systematic comparison across black-box, reasoning model variants and open-source models’ training dynamics during the alignment process, we find patterns of higher informational than social utility among reasoning models compared to their non-reasoning counterparts, as well as in open-source models that are purported to be stronger in mathematical reasoning. Our results for the open-source suite also suggest that the largest shifts in utility values happen within the first quarter of training and that the choice of base model and its pretraining data may have an outsized impact on the resulting weighting of utilities compared to the choice of feedback dataset or alignment method. Across both evaluation suites, we find that LLMs’ choice of utterances in the polite speech task is sensitive to the weightings of utilities proposed by the cognitive model. While our work studies a particular set of value trade-offs, we show that this method is responsive to diverse aspects of a rapidly evolving LLM landscape.
In providing finer-grained accounts of the mechanisms underlying high-level behavioral concepts, we propose that even behavior-specific cognitive models such as the one we consider for politeness, can be used to form and test hypotheses about other behaviors. In particular, we consider how recent concerns of sycophancy in LLMs [53, 56, 55, 17] can be described by a combination of high projected social utility, and high presentational utility, but low actual information and social utilities [cf. 11]. Throughout our results, we do not find strong examples of the described pattern among the models we test, suggesting that this may not currently a widespread safety concern. However, applying our method to known examples of sycophantic LLMs [e.g. 64] or models explicitly trained to be sycophantic [e.g. 56] could help validate such hypotheses and inform points of intervention in model training to prevent such behaviors.
Though the choices of values and goals used to construct the cognitive model in our work have been ecologically validated through human behavioral studies, they are certainly not the only goals that people entertain in communication, and further, might not be the particular set of goals that best describe LLM behaviors. Previous work has demonstrated that machine intelligence differs from our own [e.g. 75], suggesting that human and machine conceptualizations of the world likely differ as well [46]. One solution might be to develop new cognitive models of human-machine communication around neologisms that bridge human concepts and their machine counterparts to allow for a more precise understanding of LLMs as unique systems in their own right [cf. 29].
@@
While our approach offers several advantages, we also recognize the limitations of the cognitive models at the center of it. Cognitive models are often bespoke to the target domain they are crafted for, and so do not easily generalize to the open-ended nature of natural language use. Exploring how to use LLMs to map open-ended natural language data to the low-dimensional, interpretable feature space required for applying cognitive models [e.g. 41] will help to expand the settings we study with such models. Fitting cognitive models to the behavioral output of LLMs also presents several technical challenges. As with trade-offs in values, there are trade-offs between increasing the complexity and expressiveness of the cognitive model, and the explanatory power of the inferred parameter values. As such, more complex models, such as the second-order speaker model $S_{2}$ in this work, could potentially pose a challenge for making robust inferences about the critical parameters in the model. Further, we use sampling-based approximate inference, and such inference may not always be guaranteed to produce stable and unbiased results under limited computing resources in practice. We see these challenges as highlighting the importance of ongoing research at the intersection of statistics and machine learning [22, 77].

### v3

@@ 4.1 Experimental vignettes
We provide models with the same set of vignettes given to human participants in Yoon et al. (2020), which describe socially sensitive situations in which a speaker must convey their judgement of a listener’s creation (e.g. a poem, presentation, cake, etc.). The speaker’s actual opinion, or true state $s$, is expressed on a scale from 1 to 5 stars, where 1 is the lowest or most negative opinion, and 5 is the highest. [^1] We deviate from the original paper’s 0-3 heart scale to provide LLMs with a scale that is most natural to their training data, particularly online reviews. We find that this 1-5 star scale captures the semantic range of the available utterance options better than the original 0-3 scale. We present models with the set of eight utterance options $u$ (four descriptor words and their negations) in a multiple choice format:
@@
Manipulations
We extend the original third-person framing of the above scenario (simulating an LLM-as-judge) to also evaluate the LLM-as-agent and LLM-as-assistant perspectives via the first- and second-person framings of these vignettes, respectively. Finally, in addition to studying model’s default utility patterns, for the closed-source model suite, we study the effects of manipulating its communicative “goals” via system and instruction prompts that instruct the model to be: informative, rather than make someone feel good; make someone feel good (social), rather than give informative feedback; and balance both (see Section B.4 and Section B.5 for modified prompts).
@@ 4.2 Inferring cognitive model parameters
We implemented the inference model in Stan (Carpenter et al., 2017), a probabilistic programming language, and used the default Hamiltonian Monte Carlo implemented in Stan (No-U-Turn sampler, Hoffman et al. (2014)) to perform approximate inference of model parameters (see Section C.2 for further implementation details).
@@
To infer our desired cognitive model parameters $\boldsymbol{\omega}$ and $\phi$, we require an estimate of the parameter $\theta$, the probability that the utterance $u$ is true of state $s$. To obtain this, we query LLMs with a modified version of the main task where the following question is appended to the above Scenario, in its original third-person framing (see Section D.2 for an example of LLMs’ responses on this sub-task):
@@ 5 Results
Figure 3: Closed-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$, and projected mixture of informational and social utilities $\phi$, according to the cognitive model for LLMs with varying degrees of reasoning budget. Dotted lines plot model-specific results under goal conditions, averaged over framings. Solid lines show mean results across models. We find that reasoning variants prioritize information-utility over social-utility, and that goal-condition prompt manipulations shift these utility patterns in predictable ways.
@@ 5.1 Human baseline
In the original study (Yoon et al., 2020), human participants were asked to assume the role of the speaker, and to choose an utterance according to one of three goal conditions: trying to be informative, trying to be social (i.e. kind), or both. The work finds that speakers who have the conflicting goals of being both informative and kind will use more indirect speech when describing a bad state (e.g. they describe a cake that deserves only 1 star as ‘not amazing’). This behavior serves to “save face” (i.e. optimize presentational and social utilities), while still conveying useful information about the true state. It suggests that humans do not eschew one of their goals to increase utility along a single dimension, but rather, choose the utterances that will jointly maximize their competing utilities.
The hatched bar group in Figure 4 shows the maximum a posteriori (MAP) estimates of the $\phi$ and $\boldsymbol{\omega}$ parameters of the $S_{2}$ model fit to human data in (Yoon et al., 2020). Here, human speakers in the ‘informative’ goal condition project a balanced, but more information-leaning weighting of information and social utilities ($\phi=$0.49) than those in the social goal or combined goal conditions (0.37 and 0.36, respectively). The relative weightings of information and social utility in $S_{2}$, $\omega_{\text{inf}}$ and $\omega_{\text{soc}}$, track with these goal conditions, while humans’ $\omega_{\text{pre}}$, their value for communicating their $\phi$ to a listener, is highest for the informative goal condition (0.62), followed by the combined condition (0.54), and finally the social condition (0.44). The relative parameter values in each goal condition provide baselines against which we can interpret a model’s value trade-offs as a result of being prompted with the same communicative goals, relative to their default (non-goal-conditioned response, dashed line).
Figure 4: Communicative goals. Comparison of the inferred weightings of informational, social, and presentational utilities, as well as the projected trade-off $\phi$ between informational and social goals, across humans and closed-source LLMs under various manipulations of the speaker’s goals. Human results were taken from Yoon et al. (2020). Error bars indicate 95% high density region averaged over three framing manipulations crossing with three levels of reasoning budgets. We find that goal-condition prompts shift LLMs’ behavior consistently across model families, but more severely than when humans are asked to take on these same goals.
@@ 5.2 Closed-source model suite
Figure 3 shows the results of fitting the responses of Anthropic, Gemini, and OpenAI’s language models across three reasoning budgets (none, low, and medium) to the second-order speaker model. We establish that our inferred parameter values generalize to a held-out test split of the closed-source model data that includes all combinations of reasoning budget, vignette framing, communicative goal, and model family via a posterior predictive check, and find that the average MSE from our inferred parameters is significantly lower than that of randomly sampled parameter values from a reasonable prior ($\mu_{\textrm{inferred}}=0.03$, $\mu_{\textrm{random}}=0.06$; z=12.49z=-12.49; $p<0.001$; see Appendix Figure 6).
We begin with analyzing models’ default behavior in the absence of any explicit communicative goals (black lines). The parameter values of $\phi$ for the first-order speaker in $S_{2}$ measures the relative mixture of informativeness and social utility that a speaker $S_{2}$ wishes the other person to infer about them. We find that across model families, reasoning variants display higher $\phi$ values—a higher projected informational utility than social utility—than their non-reasoning counterparts. A linear mixed-effects model predicting the posterior mean $\phi$ from degrees of reasoning effect222model formula: phi $\sim$ reasoning_effort + (1|llm_family) + (1|framing) (reference level: no reasoning) with random intercepts of model family and vignette framing suggested a significant effect of both low and medium reasoning effort compared to the no-reasoning counterpart for default model behaviors ($\beta_{\text{low}}=0.228,t=6.338,p<.001$; $\beta_{\text{medium}}=0.211,t=5.846,p<.001$). The difference of the inferred $\phi$ among models of low and medium reasoning effort was not significant ($p=0.627$).
@@
These patterns of higher informational utility are similar to those seen in the inferred parameter values of $\boldsymbol{\omega}$, measure the weightings of informational, social, and presentational utilities used by the second-order pragmatic speaker. Within the Anthropic model family, Claude-Sonnet-3.5 (no reasoning), shows a significantly lower weighting of informational utility $\omega_{\text{inf}}$ compared to its low-reasoning counterpart, Claude-Sonnet-3.7 (t=5.49,p=0.005t=-5.49,p=0.005), but significantly higher social utility $\omega_{\text{soc}}$ ($t=7.17,p=0.001$). Among the OpenAI models, a similar pattern holds with significantly lower $\omega_{\text{inf}}$ for no reasoning compared to low reasoning effort (t=5.07,p=0.007t=-5.07,p=0.007), but not $\omega_{\text{soc}}$ ($p=0.06$). Conversely, the Gemini-Flash models do not show a significant difference between reasoning and non-reasoning variants for any of $\boldsymbol{\omega}$ ($p=0.43$ for $\omega_{\text{inf}}$, $p=0.20$ for $\omega_{\text{soc}}$, $p=0.89$ for $\omega_{\text{pre}}$).
@@
Finally, considering the mean speaker optimality $\alpha$, averaged over reasoning variants and vignette framings, suggests that the above described weightings of utilities do factor into the models’ choice of utterances, with all model families’ $\alpha$ being higher than 1 ($\alpha_{\text{Anthropic}}$=3.52 [3.13, 3.89]; $\alpha_{\text{Gemini}}$=6.19 [5.50, 6.88]; $\alpha_{\text{OpenAI}}$=4.78 [3.93, 5.65]).
5.2.1 Manipulating communicative goals
We find that prompting models to assume particular communicative goals shifts their behavior in interpretable ways that are consistent across model families. However, these effects of simulating these goals are much more pronounced for models than for humans. For example, in models, the informative goal condition (blue) is clearly interpreted through high $\omega_{\text{inf}}$ values, especially for the GPT models, and by a sharp increase in $\phi$ relative to both humans and the models’ own default behavior. The social goal condition is surfaced primarily through an increase in $\omega_{\text{pre}}$, a decrease in $\omega_{\text{inf}}$, and a sharp reduction (social utility-leaning) in $\phi$, across all model families. Via the $\phi$ parameter, we also see that when prompted to take on both goals, all model families project a more balanced mixture of social and informative goals relative to the highly informational default. Finally, the relative values of $\omega_{\text{inf}}$ for all the goal conditions closely resemble the human signature, while the patterns of $\omega_{\text{pre}}$ and $\omega_{\text{soc}}$ do not, suggesting that informational utility might be more easily-captured and stably-represented in models.
5.2.2 Signatures of sycophancy
In providing finer-grained accounts of the mechanisms underlying high-level behavioral concepts, we propose that even behavior-specific cognitive models such as the one we consider for politeness, can be used to form and test hypotheses about other behaviors. In particular, we hypothesized that recent concerns of sycophancy in LLMs (Liu et al., 2025; Marks et al., 2025; Malmqvist, 2024; Fanous et al., 2025) could be described by a combination of high projected social utility via a low $\phi$ and high presentational utility $\omega_{\text{pre}}$, but low actual information $\omega_{\text{inf}}$ and social $\omega_{\text{soc}}$ utilities (cf. Cheng et al., 2025). In Figure 3, we observe exactly this pattern in the social goal condition (red line), where models were prompted to act as “an assistant that wants to make someone feel good, rather than give informative feedback.” Compared to their default behavior, in this goal condition, all model families converged to such “sycophantic” utility values, with the sharpest changes occurring at the transition from no reasoning to a low reasoning budget. This reinforcement of the attributes of the system prompt does not appear as pronounced for the informative (blue) or both (purple) goal-conditions, suggesting that the content of reasoning traces may more strongly reinforce certain behavioral attributes compared to others. Applying our method to models explicitly trained to be sycophantic (e.g. Marks et al., 2025) could help further validate these findings and inform points of intervention in model training to prevent such behaviors.
Figure 5: Open-source LLM results. Inferred values of informational, social, and presentational utilities $\boldsymbol{\omega}$ (purple), and projected mixture of informational and social utilities $\phi$ (magenta), according to a cognitive model for LLMs’ training checkpoints across the RLHF process. Line variants indicate different combinations of base model and feedback dataset; rows = alignment method. Error bars indicate 95% high density region averaged across results from three framing manipulations. We find the largest shifts in values within the first quarter of training, with persistent effects of the choice of base model and pretraining data, compared to feedback dataset or alignment method.
@@ 5.3 Open-source model suite
Figure 5 shows the training dynamics of two base open-source LLMs, Qwen2.5-7B-Instruct (lighter) and Llama-3.1-8B-Instruct (darker), aligned to the UltraFeedback (dashed line) and Anthropic HH-RLHF (solid line) datasets, via DPO (top row) and PPO (bottom row). Across the different inferred parameters, we observe a number of consistent patterns within combinations of model and dataset. Across both PPO and DPO and the two feedback datasets, Qwen-instruct shows a higher $\omega_{\text{inf}}$, but lower weighting of $\omega_{\text{pre}}$ than Llama-instruct. The differences between the models’ weighting of social utility $\omega_{\text{soc}}$ are less pronounced, but still present, with Qwen-instruct generally converging to a lower weighting of social utility than Llama-instruct. The projected weighting towards informational utility in Qwen-instruct’s $\phi$, as well as its higher $\omega_{\text{inf}}$ compared to Llama-instruct aligns with prior work highlighting Qwen’s superior performance in mathematical and reasoning tasks compared to Llama (Gandhi et al., 2025; Zeng et al., 2025).
Turning to the effects of feedback dataset, we find that alignment to the UltraFeedback dataset most clearly results in convergence to a higher $\omega_{\text{inf}}$ for both base LLMs, than when aligned to Anthropic’s HH-RLHF dataset. In the case of $\omega_{\text{soc}}$, these differences are more pronounced as a result of DPO alignment, but still visible in the PPO case: for both base LLMs, alignment to HH-RLHF appears to result in a higher weighting of social utility than alignment to UltraFeedback. This aligns with the stated characteristics and attributes of the respective datasets, where HH-RLHF is a human feedback dataset that emphasizes more prosocial qualities like harmlessness and helpfulness, whereas UltraFeedback is a synthetic feedback dataset that contains more diverse instruction-following preferences.
@@
In general, we see that the largest shifts in utility values across all four parameters happen within the first quarter of training, consistent with earlier findings on rapid adaptation during RL post-training in mathematical domains (Zhao et al., 2025). While such prior work has emphasized the significance of the base model and its pretraining data Itzhak et al. (2025), our use of a shared supervised fine-tuning (SFT) stage on the same preference datasets across all models may attenuate these differences. Moreover, the relatively minor distinction between PPO and DPO in our results may be partly due to training both methods for only a single epoch, and the fact that the Armo-RM reward model used in PPO, was trained on subsets of the same UltraFeedback and Anthropic HH-RLHF datasets, further reducing divergence between the two approaches.
@@ 6 Discussion
While our approach offers several advantages, we also recognize the limitations of the cognitive models at the center of it. Cognitive models are often bespoke to the target domain they are crafted for, and so do not easily generalize to the open-ended nature of natural language use. Exploring how to use LLMs to map open-ended natural language data to the low-dimensional, interpretable feature space required for applying cognitive models (e.g. Jian & N, 2024; Qiu et al., 2025) will help to expand the settings we study with such models. Our initial attempts to do so revealed technical challenges at a number of points, first in getting smaller models to generate an alternative utterance set that captures the full semantic range of the true states we model, and later when trying to use the continuation probabilities of these generated utterances within the RSA model (c.f. Hu & Levy, 2023).
@@
Fitting cognitive models to the behavioral output of LLMs also presents several technical challenges. More complex models, such as the second-order speaker model $S_{2}$ in this work, could potentially pose a challenge for making robust inferences about the critical parameters in the model. Further, we use sampling-based approximate inference, and such inference may not always be guaranteed to produce stable and unbiased results under limited computing resources in practice. These challenges highlight the importance of ongoing research at the intersection of statistics and machine learning (Gelman et al., 2021; Shen & Broderick, 2025).
Though the choices of values and goals used to construct the cognitive model in our work have been ecologically validated through human behavioral studies, they are certainly not the only goals that people entertain in communication, and further, might not be the particular set of goals that best describe LLM behaviors. Previous work has demonstrated that machine intelligence differs from our own (e.g. Schut et al., 2025), suggesting that human and machine conceptualizations of the world likely differ as well (Kim, 2022). One solution might be to develop new cognitive models of human-machine communication around neologisms that bridge human concepts and their machine counterparts to allow for a more precise understanding of LLMs as unique systems in their own right (cf. Hewitt et al., 2025).
Lastly, we hope that such theory-driven methods of interfacing with LLMs could also be of benefit to cognitive science. These large-scale learning systems could serve as a testbed for exploring how human social intelligence may have evolved, by allowing us to probe which trade-offs emerge naturally as a result of different architectures or training decisions, and which require explicit shaping(e.g. Bonawitz et al., 2019).